OpenStudy (anonymous):

-12/20(9/18) need in simplest form

OpenStudy (anonymous):

multiply and then simplify

OpenStudy (anonymous):

just multiply -12 * 9 = -108 20 * 18 = 360 -108/360 = -.3 = -3/10